# Hydrodynamical modeling of SN 2025kg associated with the Fast X-ray Transient EP250108a

### v1

@@
Supernovae (SNe) associated with X-Ray Flashes (XRFs) are extremely rare. Therefore, the discovery of each new object in this class offers a unique opportunity to improve our understanding about their origins and potential connection with other high-energy phenomena. SN 2025kg is one of the most recent events discovered in this category, and exhibits a double-peaked light curve, with an initial cooling phase followed by the main peak. Here, we investigate the possible mechanisms powering its bolometric light curve and expansion velocities, using numerical calculations to simulate the explosion. We found that low ejecta masses (M M⊙) and moderate explosion energies (E erg) are required to reproduce the data. Our models also show that a large amount of nickel (M) is needed to achieve the high luminosity of SN 2025kg, which makes this scenario difficult to sustain. As an alternative, we explore a model in which a millisecond magnetar serves as the primary energy source. A magnetar with a spin period of ms and a magnetic field of 28 G give an adequate match to the data. To account for the early cooling phase, we assume the presence of a dense circumstellar material surrounding the progenitor, with a mass of 0.27 M⊙ and an extension of 500 R. A comparison and modeling of a select group of SNe –SN 2006aj, SN 2020bvc and SN 2023pel– is also presented. A remarkable similarity emerges between SN 2025kg and SN 2023pel. As SN 2023pel was recently proposed to be powered by a magnetar, this further supports the magnetar scenario for SN 2025kg.
@@
The connection between long gamma-ray bursts (GRB) and core-collapse supernovae (CCSNe) is now well established (see Cano et al. (2017) for a review). Numerous associations have now been confirmed, with all corresponding to SNe classified as hydrogen- and helium-deficient objects, with broad spectral lines indicative of large kinetic energies (SNe Ic-BL). Although there are several confirmations, these explosions are still very rare compared to other types of CCSNe. Even less frequent are those SNe accompanying X-ray Flashes (XRFs111Also known as Fast X-ray Transients, FXTs.). The first clear identification of these events was SN 2006aj (Soderberg et al. 2006; Pian et al. 2006), followed a few years later by SN 2010bh (Cano et al. 2011; Olivares et al. 2012). Since then, no further associations had been reported until very recently, with the launch of the Einstein Probe mission (EP, Yuan (2022)). Given the scarcity of these events, any new object of this type deserves to be studied in detail, as it may help to better understand the connections between XRF-SNe, GRB-SNe, and SNe Ic-BL.
Recently, a new XRF-SN association has been reported (Li et al. 2025). The FXT (EP250108a) was discovered by the EP mission on 8 January 2025, and the following observations confirmed the existence of an optical counterpart, designated as SN 2025kg and classified as SN Ic-BL (Eyles-Ferris et al. 2025). Since this discovery, several works have been published presenting and analyzing the photometric and spectroscopic properties of this event Rastinejad et al. (2025) (R25), Li et al. (2025) (L25), Srinivasaragavan et al. (2025) (S25), and Eyles-Ferris et al. (2025) (EF25).
@@
In previous works, the properties of SN 2025kg were analyzed using an analytical prescription. Here, we modeled the LC and expansion velocities using numerical simulations, based on public data (see § 2). Our results are presented in § 3. Comparing SN 2025kg with a subset of SNe associated with high-energy radiation, we find a striking resemblance to SN 2023pel –the most recent GRB-SN association– which was recently modeled by us, assuming a magnetar as the main power source (Roman Aguilar et al. (2025, hereafter RA25)). A detailed comparison with previous studies is provided in §4, and the main conclusions are summarized in § 5.
@@ 2 Models and Observations
The bolometric LCs of SNe are very sensitive to both the physical parameters of their progenitor stars and the underlying power sources. Conversely, the early evolution depends strongly on the possible presence of nearby CSM. It is common to compare theoretical LC models with observations in order to derive the progenitor and explosion parameters. However, an important degeneracy exists between parameters when only photometric data is used. Including the photospheric velocity evolution, as inferred from some spectral lines, can help to break this degeneracy. Fe II velocities have been proposed as an effective tracer of photospheric velocity (Dessart & Hillier 2005). Here, we used Fe II velocity measures by S25 to compare with our models.
Various authors have estimated the pseudo- or bolometric luminosity of SN 2025kg, and their results are overall in remarkable agreement (see black symbols in Fig. 1). S25 produced a bolometric LC by applying the color-calibration method provided by Lyman et al. (2014); their data covered the early cooling phase and the main peak from 1 to 35 days. L25 computed a pseudo-bolometric LC through direct flux integration over rest-frame wavelengths of 3000-9000. The data provided by R25 was calculated using grizJHK photometry and co-bands from ATLAS. Although the three LCs are similar around the main peak, significant discrepancies appear at early epochs (t ¡ 6 days) during the first peak (cooling phase). It is possible that the non-inclusion of a UV correction accounts for these differences, even though such a correction carries high uncertainties. Accordingly, the earliest data–and parameters derived from them–should be treated with caution.
@@
As an initial input for our simulation, a pre-SN model representing the state of the star prior to the explosion is required. Hydrostatic structures, calculated with stellar evolution codes, are typically employed as pre-SN conditions. However, the outermost regions of these structures are often modified by hand when the effect of a CSM is incorporated, since no evolutionary models consistently reproduce the CSM conditions required to match the observations. In this work, we use the same grid of stellar models recently employed in RA25. These models consist of H-free structures of different masses, specifically stars with main sequence mass (MZAMS) of 13, 15, 18, 20, and 25 M⊙ which correspond to pre-SN mass of 3.3, 4, 5, 6, and 8 M⊙. All these models have a compact structure at the explosion time with R 5 before modifying its structure, considering the inclusion of CSM.
@@ 3 Results
Figure 1 shows a comparison of SN 2025kg with other energetic SNe. SN 2006aj associated with an XRF (Šimon et al. 2010; Pian et al. 2006), SN 2020bvc, a SN Ic-BL with some evidence of an offset jet (Ho et al. 2020), and SN 2023pel, the most recent GRB-SN association (Srinivasaragavan et al. (2024); Hussenot-Desenonges et al. (2024)). From the figure, we highlight some aspects: (1) there is a clear diversity in luminosity within the XRF-SN objects, (2) early emission prior to the main peak is present in all the comparison events. Interestingly, this early emission appears to be more frequent in SNe associated with high-energy radiation than in normal SN Ic, and it may provide some insights into their progenitor origin. (3) Despite the large range of radiative output (luminosity), the kinetic energy (velocities) seems to exhibit a more homogeneous behavior, and (4) SN 2025kg is remarkably similar to SN 2023pel in terms of both luminosity and the Fe II velocities.
@@
Figure 1 presents some of our preferred models for SN 2025kg. A model powered only by 56Ni (the Ni model) is shown in a black dashed line, while a magnetar plus some 56Ni (the magnetar model, for simplicity) is shown in a black dash-dotted line. Some amount of nickel ( 0.2 M⊙) was also included in the magnetar model to improve the agreement with the observations at later times (t days). Also, note that some 56Ni is naturally expected to be synthesized during the explosion. Finally, the black solid line represents our preferred model, consisting of a magnetar plus a CSM (Mag+CSM) model. The parameters of each of these models are listed in Table 1.
All the models presented in Figure 1 reproduce the data of SN 2025kg reasonably well around the main peak. Note that models that include a magnetar contribution require a more massive progenitor (5 M⊙; see Tab. 1) than the Ni model. The same was previously noted in our analysis of SN 2023pel. This is because the magnetar model supplies additional energy, which results in a narrower LC. To counterbalance this effect and reproduce the observed LC width, a more massive progenitor is needed to achieve the necessary broadening. The Fe II velocities are also well reproduced, except for the first data point, which none of our models could reproduce. It is possible that the measurement of this line velocity at early epochs is subject to large uncertainties.222Alternatively, a different CSM distribution than a steady wind could help to improve the agreement (see the inset in Fig. 1).. Here, we present only our favorite models, although several alternatives were previously explored to select an acceptable solution in each scenario. These calculations involved different progenitor masses and variations in the free model parameters (E, MNi, and P and B). However, given the similarity between SN 2025kg and SN 2023pel, the exploration was guided by our previous results.
@@
In this work, we only explore the possibility that CSM interaction is responsible for powering the early emission. To include the effect of the CSM, we attached some material in the outermost layer of the pre-SN density profile assuming a stationary wind law (). After exploring various configurations, primarily adjusting the CSM extension and mass, we identified a model that accurately reflects the data. This model is depicted in Fig. 1 with a black solid line and has a CSM extension and mass of 500 R⊙ and 0.27 M⊙, respectively. A wind velocity of 115 was assumed in our calculations. Although CSM models reproduce the initial LC well, they yield a poorer match to the first velocity data. In fact, our CSM models produce even lower velocities at those times. However, adopting a different CSM distribution–rather than the steady wind used here–could improve the velocity match without significantly affecting the initial LC behavior (see the inset in Fig. 1).
In Fig. 1 we also present models for the SN 2006aj (magenta) and SN 2020bvc (cyan) calculated using the same code and progenitor grid. The parameters of the models are given in Tab. 1. In these cases, only a nickel power source was explored, due to the relatively normal luminosities. For modeling the early phase, we attached the CSM on top of the external density profile, as explained above for SN 2025kg. The values found for MNi are within the expected range for H-free SNe and are considerably lower than those found for SN 2025kg (Ni model), in agreement with their maximum luminosities. On the other hand, SN 2020bvc has the highest E of the sample, consistent with the behavior of its velocities, whose values remain higher throughout their evolution. SN 2006aj shows the lower MZAMS and Mej, in agreement with showing the narrowest LC. The properties of the CSM are also in concordance with the behavior shown in their LC, as a slower decay in the early phase is associated with higher values of MCSM and RCSM333Although some degeneration between MCSM and RCSM exists., which in turn produces higher values of the LC minimum. Finally, we note that the model parameters of SN 2023pel are in very good agreement with the value obtained for our preferred model for SN 2025kg (Mag+CSM) which is expected given the similarities between both SNe.
@@ 5 Conclusions
Our numerical models indicate that a large amount of nickel (M 0.85 M⊙) and a low ejecta mass (M 1.9 M⊙) are required to explain the observations when considering a model powered only by 56Ni. Alternatively, a model including an additional energy source provided by a magnetar—with P = 2.9 ms and B = G—and a typical nickel mass (0.2 M⊙), can also reproduce the observations. As in the case of SN 2023pel, we also favor the magnetar scenario for SN 2025kg, as we think this model is physically more plausible (see RA25). On the other hand, the early LC component was modeled by assuming the presence of some CSM located near the progenitor star before the explosion. By adopting an extension of 500 R⊙ and a CSM mass of 0.27 M⊙, we were able to reproduce the early-time emission under the assumption of a steady wind profile.
When comparing the properties of SN 2025kg with those of other events associated with high-energy radiation, we find that their luminosities are highly diverse, while their expansion velocities exhibit much smaller variations. Conversely, an early emission component appears to be relatively common among these objects, unlike in most normal He-deficient SNe.

### v2

@@
Supernovae (SNe) associated with X-Ray flashes (XRFs) are extremely rare. Therefore, the discovery of each new object in this class offers a unique opportunity to improve our understanding of their origins and potential connection with other high-energy phenomena. SN 2025kg is one of the most recent events discovered in this category and exhibits a double-peaked light curve, with an initial cooling phase followed by the main peak. Here, we investigate the possible mechanisms powering its bolometric light curve and expansion velocities, using numerical calculations to simulate the explosion. We find that low ejecta masses (M M⊙) and moderate explosion energies (E erg) are required to reproduce the data. Our models also show that a large amount of nickel (M) is needed to achieve the high luminosity of SN 2025kg, which makes this scenario difficult to sustain. As an alternative, we explore a model in which a millisecond magnetar serves as the primary energy source. A magnetar with a spin period of ms and a magnetic field of 28 G provides an adequate match to the data. To account for the early cooling phase, we assume the presence of a dense circumstellar material surrounding the progenitor, with a mass of 0.27 M⊙ and an extension of 500 R⊙. A comparison and modeling of a selected group of SNe—SN 2006aj, SN 2020bvc, and SN 2023pel—is also presented. A remarkable similarity emerges between SN 2025kg and SN 2023pel. Since SN 2023pel was recently proposed to be powered by a magnetar, this further supports the magnetar scenario for SN 2025kg.
@@
The connection between long gamma-ray bursts (GRBs) and core-collapse supernovae (CCSNe) is now well established (see Cano et al. (2017) for a review). Numerous associations have been confirmed, with all corresponding to SNe classified as hydrogen- and helium-deficient objects, with broad spectral lines indicative of large kinetic energies (SNe Ic-BL). Although there are several confirmations, these explosions are still very rare compared to other types of CCSNe. Even less frequent are those SNe accompanying X-Ray flashes (XRFs111Also known as Fast X-ray Transients, FXTs.). The first clear identification of these events was SN 2006aj (Soderberg et al. 2006; Pian et al. 2006), followed a few years later by SN 2010bh (Cano et al. 2011; Olivares et al. 2012). Since then, no further associations have been reported until very recently. Given the scarcity of these events, every new object of this type deserves detailed study, as it may help to better understand the connections between XRF-SNe, GRB-SNe, and SNe Ic-BL.
Recently, a new XRF-SN association has been reported (Li et al. 2025). The FXT (EP250108a) was discovered by the EP mission on 8 January 2025, and subsequent observations confirmed the existence of an optical counterpart, designated as SN 2025kg and classified as SN Ic-BL (Eyles-Ferris et al. 2025). Since this discovery, several works have been published presenting and analyzing the photometric and spectroscopic properties of this event Rastinejad et al. (2025) (R25), Li et al. (2025) (L25), Srinivasaragavan et al. (2025) (S25), and Eyles-Ferris et al. (2025) (EF25).
@@ 2 Models and observations
The bolometric LCs of SNe are very sensitive to their progenitor properties and power sources, while their early evolution can be strongly influenced by a nearby CSM. It is common to compare theoretical LC models with observations in order to derive the progenitor and explosion parameters. However, an important degeneracy arises when only photometric data are used. Including the photospheric velocity evolution, as inferred from some spectral lines, can help to break this degeneracy. Fe II velocities have been proposed as an effective tracer of photospheric velocity (Dessart & Hillier 2005). Here, we used Fe II velocity measures by S25 to compare with our models.
Various authors have estimated the pseudo- or bolometric luminosity of SN 2025kg, and their results are overall in remarkable agreement (see black symbols in Fig. 1). S25 produced a bolometric LC by applying the color-calibration method provided by Lyman et al. (2014); their data covered the early cooling phase and the main peak from 1 to 35 days. L25 computed a pseudo-bolometric LC through direct flux integration over rest-frame wavelengths of 3000-9000. The data provided by R25 was calculated using grizJHK photometry and co-bands from ATLAS. Although the three LCs are similar around the main peak, significant discrepancies appear at early epochs (t ¡ 6 days) during the first peak (cooling phase). It is possible that the non-inclusion of a UV correction accounts for these differences, even though such a correction carries high uncertainties. Accordingly, the earliest data—and parameters derived from them—should be treated with caution.
@@
As an initial input for our simulation, a pre-SN model representing the state of the star prior to the explosion is required. Hydrostatic structures, calculated with stellar evolution codes, are typically employed as pre-SN conditions. However, the outermost regions of these structures are often modified by hand when the effect of a CSM is incorporated, since no evolutionary models consistently reproduce the CSM conditions required to match the observations. In this work, we use the same grid of stellar models recently employed in Roman Aguilar et al. (2025, hereafter RA25). These models consist of H-free222He-free progenitors would be more appropriate for SNe Ic, but no self-consistent pre-SN models with complete He removal are currently available. structures of different masses, specifically stars with main-sequence mass (MZAMS) of 13, 15, 18, 20, and 25 M⊙ which correspond to pre-SN mass of 3.3, 4, 5, 6, and 8 M⊙. All these models have a compact structure at the explosion time with R 5 before modifying its structure, considering the inclusion of CSM.
@@ 3 Results
Figure 1 compares SN 2025kg with other energetic SNe. SN 2006aj associated with an XRF (Šimon et al. 2010; Pian et al. 2006), SN 2020bvc, a SN Ic-BL with some evidence of an offset jet (Ho et al. 2020), and SN 2023pel, the most recent GRB-SN (Srinivasaragavan et al. (2024); Hussenot-Desenonges et al. (2024)). From the figure, we highlight some aspects: (1) there is a clear diversity in luminosity within the XRF-SN objects, (2) early emission prior to the main peak is present in all the comparison events. Interestingly, this early emission appears to be more frequent in SNe associated with high-energy radiation than in normal SN Ic, and it may provide some insights into their progenitor origin. (3) Despite the large range of radiative output (luminosity), the kinetic energy (velocities) seems to exhibit a more homogeneous behavior, and (4) SN 2025kg is remarkably similar to SN 2023pel in terms of both luminosity and the Fe II velocities.
@@
Figure 1 presents some of our preferred models for SN 2025kg; a model powered only by 56Ni (the Ni model) and a magnetar plus some 56Ni (the magnetar model, for simplicity). About 0.2 M⊙ of nickel was also included in the latter to improve the agreement with the observations at later times (t days), and because some 56Ni is naturally expected to be synthesized during the explosion. Finally, our selected model consists of a magnetar plus a CSM (Mag+CSM). The parameters of each of these models are listed in Table 3.
All the models presented in Fig. 1 reproduce the data of SN 2025kg reasonably well around the main peak. Note that models including a magnetar contribution require a more massive progenitor (5 M⊙; see Tab. 3) than the Ni model. The same was previously noted in our analysis of SN 2023pel. This is because the magnetar model supplies additional energy, which results in a narrower LC. To counterbalance this effect and reproduce the observed LC width, a more massive progenitor is needed to achieve the necessary broadening. The Fe II velocities are also well reproduced, except for the first data point, which none of our models could reproduce. It is possible that the measurement of this line velocity at early epochs is subject to large uncertainties.444Alternatively, a different CSM distribution than a steady wind could help to improve the agreement (see the inset in Fig. 1).. Here, we present only our favorite models, although several alternatives were previously explored to select an acceptable solution in each scenario. These calculations involved different progenitor masses and variations in the free model parameters (E, MNi, and P and B). However, given the similarity between SN 2025kg and SN 2023pel, the exploration was guided by our previous results.
@@
In this work, we only explore the possibility that CSM interaction powers the early emission. To include the effect of the CSM, we attached some material in the outermost layer of the pre-SN density profile, assuming a stationary wind law (). After exploring various configurations, primarily adjusting the CSM extension and mass, we identified a model that accurately reflects the data. This model is presented in Fig. 1 and has a CSM extension and mass of 500 R⊙ and 0.27 M⊙, respectively. A wind velocity of 115 was assumed in our calculations. Although CSM models reproduce the initial LC well, they yield a poorer match to the first velocity data. In fact, our CSM models produce even lower velocities at those times. However, adopting a different CSM distribution–rather than the steady wind used here–could improve the velocity match without significantly affecting the initial LC behavior (see the inset in Fig. 1).
In Fig. 1 we also present models for the SN 2006aj and SN 2020bvc calculated using the same code and progenitor grid. The parameters of the models are given in Tab. 3. In these cases, only a nickel power source was explored, due to the relatively normal luminosities. For modeling the early phase, we attached the CSM on top of the external density profile, as explained above for SN 2025kg. The values found for MNi are within the expected range for H-free SNe and are considerably lower than those found for SN 2025kg (Ni model), in agreement with their maximum luminosities. On the other hand, SN 2020bvc has the highest E of the sample, consistent with the behavior of its velocities, whose values remain higher throughout their evolution. SN 2006aj shows the lower MZAMS and Mej, in agreement with showing the narrowest LC. The properties of the CSM are also in concordance with the behavior shown in their LC, as a slower decay in the early phase is associated with higher values of MCSM and RCSM555Although some degeneration between MCSM and RCSM exists., which in turn produces higher values of the LC minimum. Finally, we note that the model parameters of SN 2023pel are in very good agreement with the value obtained for our preferred model for SN 2025kg (Mag+CSM) which is expected given the similarities between both SNe.
@@ 5 Conclusions
Our numerical models indicate that a large amount of nickel (M 0.85 M⊙) and a low ejecta mass (M 1.9 M⊙) are required to explain the observations when considering a model powered only by 56Ni. Alternatively, a model including an additional energy source provided by a magnetar—with P = 2.9 ms and B = G—and a typical nickel mass (0.2 M⊙), can also reproduce the observations. As in the case of SN 2023pel, we also favor the magnetar scenario for SN 2025kg, as we think this model is physically more plausible (see RA25). On the other hand, the early LC component was modeled by assuming the presence of some CSM located near the progenitor star before the explosion. By adopting an extension of 500 R⊙ and a CSM mass of 0.27 M⊙, we were able to reproduce the early-time emission under the assumption of a steady wind profile. These parameters imply an extreme mass-loss episode 0.1 yr before the explosion, which may be difficult to justify physically. An alternative is that the progenitor had an extended envelope, as that needed to reproduce early data in SN IIb (Bersten et al. 2012). Such an envelope could naturally arise from binary mass transfer, which could also favor the magnetar formation scenario.
When comparing the properties of SN 2025kg with those of other events associated with high-energy radiation, we find that their luminosities are highly diverse, while their expansion velocities exhibit much smaller variations. Conversely, an early emission component appears to be relatively common among these objects, unlike in most normal He-deficient SNe. Some of the observed characteristics seem easier to explain within a binary evolution scenario. We therefore speculate that binary progenitors are likely responsible for this class of high-energy transients.
@@ Acknowledgements.
M.B. acknowledges support from PIP 112-202001-10034 and PICT 2020-1141 grants and to CSIC-iCOOP exchange program.References
- Bersten et al. (2011) Bersten, M. C., Benvenuto, O., & Hamuy, M. 2011, ApJ, 729, 61
- Bersten et al. (2012) Bersten, M. C., Benvenuto, O. G., Nomoto, K., et al. 2012, ApJ, 757, 31
- Bersten et al. (2016) Bersten, M. C., Benvenuto, O. G., Orellana, M., & Nomoto, K. 2016, ApJ, 817, L8
- Bersten et al. (2024) Bersten, M. C., Orellana, M., Folatelli, G., et al. 2024, A&A, 681, L18
- Cano et al. (2011) Cano, Z., Bersier, D., Guidorzi, C., et al. 2011, ApJ, 740, 41
- Cano et al. (2017) Cano, Z., Wang, S.-Q., Dai, Z.-G., & Wu, X.-F. 2017, Advances in Astronomy, 2017, 8929054
- Dessart & Hillier (2005) Dessart, L. & Hillier, D. J. 2005, A&A, 439, 671
- Englert Urrutia et al. (2020) Englert Urrutia, B. N., Bersten, M. C., & Cidale, L. S. 2020, BAAA, 61B, 51
- Ertini et al. (2025) Ertini, K., Regna, T. A., Ferrari, L., et al. 2025, A&A, 699, A60
- Eyles-Ferris et al. (2025) Eyles-Ferris, R. A. J., Jonker, P. G., Levan, A. J., et al. 2025, ApJ, 988, L14
- Gutiérrez et al. (2021) Gutiérrez, C. P., Bersten, M. C., Orellana, M., et al. 2021, MNRAS, 504, 4907
- Ho et al. (2020) Ho, A. Y. Q., Kulkarni, S. R., Perley, D. A., et al. 2020, ApJ, 902, 86
- Hussenot-Desenonges et al. (2024) Hussenot-Desenonges, T., Wouters, T., Guessoum, N., et al. 2024, MNRAS, 530, 1
- Li et al. (2025) Li, W. X., Zhu, Z. P., Zou, X. Z., et al. 2025, arXiv e-prints, arXiv:2504.17034
- Lyman et al. (2014) Lyman, J. D., Bersier, D., & James, P. A. 2014, MNRAS, 437, 3848
- Martinez et al. (2022) Martinez, L., Bersten, M. C., Anderson, J. P., et al. 2022, A&A, 660, A41
- Olivares et al. (2012) Olivares, E. F., Greiner, J., Schady, P., et al. 2012, in IAU Symposium, Vol. 279, Death of Massive Stars: Supernovae and Gamma-Ray Bursts, 375–376
- Orellana & Bersten (2022) Orellana, M. & Bersten, M. C. 2022, A&A, 667, A92
- Orellana et al. (2018) Orellana, M., Bersten, M. C., & Moriya, T. J. 2018, A&A, 619, A145
- Pian et al. (2006) Pian, E. et al. 2006, Nature, 442, 1011
- Rastinejad et al. (2025) Rastinejad, J. C., Levan, A. J., Jonker, P. G., et al. 2025, ApJ, 988, L13
- Román Aguilar & Bersten (2023) Román Aguilar, L. M. & Bersten, M. C. 2023, BAAA, 64, 68
- Roman Aguilar et al. (2025) Roman Aguilar, L. M., Saez, M. M., Ertini, K., & Bersten, M. C. 2025, A&A, 698, A78
- Soderberg et al. (2006) Soderberg, A. M., Kulkarni, S. R., Nakar, E., et al. 2006, Nature, 442, 1014
- Srinivasaragavan et al. (2025) Srinivasaragavan, G. P., Hamidani, H., Schroeder, G., et al. 2025, ApJ, 988, L60
- Srinivasaragavan et al. (2024) Srinivasaragavan, G. P., Swain, V., O’Connor, B., et al. 2024, ApJ, 960, L18
- Sutherland & Wheeler (1984) Sutherland, P. G. & Wheeler, J. C. 1984, ApJ, 280, 282
- Taddia et al. (2018) Taddia, F., Stritzinger, M. D., Bersten, M., et al. 2018, A&A, 609, A136
- Šimon et al. (2010) Šimon, V., Pizzichini, G., & Hudec, R. 2010, A&A, 523, A56